# Revitalising the Single Batch Environment: A ’Quest’ to Achieve Fairness and Efficiency

### v2

@@ Abstract
In the realm of computer systems, efficient utilisation of the CPU (Central Processing Unit) has always been a paramount concern. Researchers and engineers have long sought ways to optimise process execution on the CPU, leading to the emergence of CPU scheduling as a field of study. In this research, we have analysed the single offline batch processing and investigated other sophisticated paradigms such as time-sharing operating systems and wildly used algorithms, and their shortcomings. Our work is been directed towards two fundamental aspects of scheduling: efficiency and fairness. We propose a novel algorithm for batch processing that operates on a preemptive model, dynamically assigning priorities based on a robust ratio, employing a dynamic time slice, and utilising periodic sorting techniques to achieve fairness. By engineering this responsive and fair model, the proposed algorithm strikes a delicate balance between efficiency and fairness, providing an optimised solution for batch scheduling while ensuring system responsiveness.
@@ 1 Introduction
00footnotetext: pre-print under review
@@
The roots of CPU scheduling [1][2] can be traced back to the early days of computing when batch processing systems were prevalent. In these systems, a set of fixed-size jobs was submitted as a batch to the computer, and the CPU had to execute them one after another. This approach suffered from inefficiencies as the CPU lacked responsiveness or processes waiting for the completion of others. To overcome certain limitations, scheduling algorithms were devised to manage the execution order of jobs, minimising idle time and maximising CPU utilisation.
@@
Efficient scheduling not only impacts system performance but also has significant economic implications. In today’s digital era, where computational power is a valuable resource, optimising CPU scheduling can lead to substantial cost savings. By ensuring that the CPU is effectively utilised and idle time is minimised, organisations can complete computational tasks faster, reduce energy consumption, and ultimately save money. In addition to efficiency, the consideration of fairness in job selection is a crucial aspect of schedulers that is frequently underestimated. However, fairness is often viewed as a subjective metric lacking a universally agreed-upon definition in various task-scheduling contexts. In this research paper, we return to the fundamental scheduling paradigm of batch processing. Here, we have a single, fixed-size queue of jobs, each with predetermined burst times. A single machine processes the queue, with preemption allowed between jobs. This paradigm forms the cornerstone of more advanced scheduling concepts, such as time-sharing, multi-programming, and even some real-time job scheduling. Despite the growing demand for advanced scheduling algorithms in recent times, this very foundational paradigm has comparatively been overlooked. As a result, the potential usage of the paradigm has also not been utilised across various domains. Hence, our research seeks to reinvigorate this single-batch environment with our novel algorithm. We have extensively analyzed classical algorithms commonly used in time-sharing and multi-programming systems [1] and their individual shortcomings, particularly in terms of both efficient and fair distribution of job selections across various probabilistic distributions of the burst times of the jobs. Through our efforts, we strive to enhance system productivity, reduce costs, and improve the overall user experience in various batch or batch-like computing environments. Across the manuscript, we’ve interchangeably used the terms ’Job’, ’Task’, and ’Process’.
1.1 CPU Scheduling
In the realm of CPU scheduling, various algorithms and techniques have been developed to address the challenges posed by process management. These algorithms consider factors such as process priority, resource requirements, scheduling policies, and system goals to determine the most suitable order of process execution. CPU scheduling algorithms range from simple heuristics to complex optimisation algorithms, each with its own advantages and limitations. These algorithms consider various factors and employ diverse strategies to optimise process execution. These algorithms aim to ensure equitable resource distribution, reduce waiting times, minimise response times, and improve overall system efficiency.
Highlights of the work
  • Integration of dynamic time slicing, preemption mechanism, and a periodic holistic sorting of processes to optimise the allocation of CPU time to processes.
  • Striking a balance between responsiveness and efficiency through thoughtful consideration of crucial process attributes and system dynamics such as removing starvation, priority inversion, unnecessary context switching, and additional computation overhead for scheduling.
  • Robust solution for single batch scheduling, offering improved system performance, fairness, and efficiency
Organisation of the paper
: The content in the proposed article is organized into several key sections. Firstly, in Section 2, we provide a comprehensive overview of existing and related works in the field. Section 3 presents our proposed work, which includes a detailed description of the algorithm and flowchart, followed by a thorough analysis. Moving on to Section 4, we present the results of our experiments and analysis and subsequently engage in a detailed discussion of our findings. In Section 5, we explore potential optimisations and improvements that can be made to the proposed algorithm. Finally, in Section 6, we offer a concise summary of the key findings and their implications for future research.
@@ 2.1 Glossary
2.2 Classical Algorithms
[1]
  • First-Come, First-Served (FCFS): The FCFS algorithm, a non-preemptive scheduling algorithm, operates on a simple principle: the process that arrives first is served first. This algorithm is easy to implement and ensures fairness in process execution. However, it suffers from poor average response time and can lead to a phenomenon known as the ”convoy effect,” where long processes hold up shorter ones.
  • Shortest Job Next (SJN) - Non-Preemptive and Preemptive Versions: The SJN algorithm, also referred to as the shortest job first algorithm, aims to minimise the average waiting time of processes. In its non-preemptive version, it selects the process with the smallest total execution time (burst time) for execution. While, it is well known that for single server systems without preemption, SJN provides optimal performance in terms of minimising waitingtime [3, 1].
  • Longest Remaining Time First (LRTF): is a preemptive CPU scheduling algorithm that selects the process with the longest remaining execution time. It allocates the CPU to the process with the greatest time requirement from the available processes. Preemption is permitted, allowing higher-priority processes with longer execution times to interrupt the currently running process.
  • Multilevel Queue and Multilevel Feedback Queue Scheduling,[2]: Multilevel Queue Scheduling categorises processes into multiple priority queues, each with its own scheduling algorithm. Processes are initially assigned to a specific queue based on their attributes, such as priority or type. This approach provides a balance between fairness and responsiveness, as different classes of processes receive appropriate treatment. However, determining the optimal number and configuration of queues can be challenging.
2.3 Recent Advancements
@@
Over time, researchers have proposed enhancements and improvements to these classical algorithms. For instance, variations of RR, such as variations of Weighted Round Robin (WRR) [4], [5] and Multilevel Queue Scheduling, have been introduced to address the limitations of strict time slicing. Additionally, policies like Priority Scheduling and Multilevel Feedback Queue Scheduling have been developed to incorporate process priority and dynamically adjust scheduling parameters. Recent advancements in CPU scheduling have focused on adaptive and intelligent algorithms, leveraging machine learning and optimisation techniques [6], [7], [8]. These algorithms aim to dynamically adapt to workload patterns, predict burst times, and improve system performance. Examples include Reinforcement Learning-based scheduling [9], [10] algorithms, fuzzy-logic based heuristics [11], [12], [13] and Genetic Algorithm-based approaches [14]. While these advanced frameworks for complex operating systems take huge additional computations, other strategies, like backfilling algorithms [15], fair-share schedulers [16], gang scheduling [17], deadline-based schedulers [18] etc are complex to implement, cost computational overhead and do not comprehend the inter-dependencies of different attributes in an easy manner and not always suitable and/or applicable to the batch processing.
The exploration of existing literature on CPU scheduling has provided valuable insights and observations that have shaped the development of our proposed algorithm. A careful analysis of prior research and real-world scenarios has highlighted the need for a scheduling approach that addresses both fairness and efficiency in the batch environment. The shortcomings of traditional algorithms, such as the trade-offs between priority and efficiency, as well as the challenges posed by process starvation and varying burst times, have driven our quest for an optimised solution.
3 Proposal of our work
@@
A batch processing system can be visualised as a fixed-size array of jobs. In this research, we have considered the case where the jobs have their predetermined bursttimes. This is the most fundamental paradigm in scheduling theory which along with its several variants serves as the cornerstone for some of the most important industrial processes as well as in various domains in data engineering, data warehousing etc. Now, our ’Quest’ is to analyse this very simple batch paradigm with the intention to investigate, primarily the behaviour of the two metrics researchers mostly interested in: Efficiency and Fairness.
@@ I. Efficiency
A batch processing system is heavily used in industrial setups [19, 20] where we are mostly concerned with the completion of all the jobs. In our work, there is no explicit deadline, penalty or priority of jobs. This is why, our primary point of interest is the average waiting and turnaround time for the whole batch for quantifying the efficiency aspect. Researchers, in the past, had also considered the number of tardy jobs [21], maximum lateness [22], and sometimes makespans [23] in a set of derivatives of the same paradigms which is not suitable and/or applicable for our current objectives of this work. Efficiency in queuing systems generally has at least one standard metric and definition to measure the performance of policies.
@@ II. Fairness
Fairness, unlike efficiency, doesn’t have a universally accepted metric to deal with [24]. Experts had mostly related fairness in terms of selection and providing proportionate timeQuantum to jobs [25, 24] in the past. However, to the best of our knowledge, there is no universal quantifiable metric of fairness for single-batch processing. In our work, we do not intend to give any fine-tuned definition or benchmark for the same but we consider ’Fairness’ to be a relative behavioural property observed over the whole batch with respect to its average turnaroundtime / waitingtime and average responsetime.
For a batch, the primary work is to have a class of algorithms that can sufficiently perform well in terms of its efficiency measure. But that is not the only primary concern for the experts to look at [25]. That’s why, practitioners often consider the average responsetime of the system along with its completion time [26], which is essentially the time duration of a job getting its first response from the time of arrival. In our setting, as all the jobs have arrived in the batch and no further addition of jobs would be there at any instance, depending on the selection of the jobs at each cycle responsetime provides an overview of how fairly jobs have been selected provided it competitively aligns its efficiency (average waitingtime / turnaroundtime ) with respect to the benchmark (SPT / SRTF). Now responsetime for policies differs fundamentally from its design perspective. For example, in RR response time is a crucial parameter but in general, it chooses timeQuantums to run the jobs with preemption if required, in an empirical manner whereas, SRTF or LRTF prioritises jobs with shorter and longer remaining bursts respectively and do not explicitly mean to be supremely effective for responsetime. For example, SJF optimally minimises the average waiting & turnaround time irrespective of the distribution the bursts follow. But the same cannot be extended for average response time.
In our work, we have emphasised the selection of the jobs at each iteration of our algorithm, discussed in subsequent sections. Unlike policies like SRTF, LRTF which greedily choose jobs based on a single parameter, we have come up with a ratio to track the changes in several important parameters based on which the scheduler holistically runs the processes. The greedy policies, although conditionally perform pretty well, are often the reason for severe starvation, and convoy effect in the system (which is not very fair).
So with that note, with respect to our work, given a distribution, fairness is essentially the behaviour, our policy exhibits due to which it refrains from choosing jobs purely greedily and chooses more holistically such that starvation and/or excessive preemption, if taking place, are reduced in the whole batch provided the policy is competitively efficient. This has been demonstrated exhaustively with respect to the comparison between average waiting and average response time across distribution in the ’Result’ section.
3.1 Proposed Work
@@
When emphatically going through examples of process clusters, a very thoughtful observation for us was to discover the fact that any cluster of processes is never discrete throughout execution. In other words, the current attributes of a particular cluster of processes will inevitably change after some interval and we would be dealing with different types of sub-problems (here cluster of processes) in each cycle. In a multilevel queue or multilevel feedback queue for example, if we have an interval of ‘t’ units and a set of algorithms ‘S’, we may halt after each ’t’ unit of time, analyse the system, and can choose the ‘most suitable’ algorithm from ‘S’. For a tiebreaker amongst comparable algorithms, we can even take their space and time complexity or other parameters into account. Multilevel queues and multilevel feedback queues (MLQs/MLFQs) are computationally intensive [1] and can be overkill when striking a balance between efficiency and fairness, this very observation was the main inspiration to come up with a metric that can address the crucial aspects and intricacies of the process attributes after a certain period. Our approach is primarily experimental and substantially empirical in nature. After experimenting with several ways of measuring the changes in various parameters and their interdependencies, we have finalised the ratio in the algorithm below based on a few aspects elaborated in 3.2 section. While this may not be as sophisticated as MLQ / MLFQ, the interdependencies of attributes, if addressed effectively, lead to a near-optimal solution. With the potential to address the limitations of traditional batch processing, our approach opens doors to more conclusive approaches in this domain.
@@
     timeQuantumnormal-$\leftarrow$ ceil((mean of remainingtime of processes + median of remainingtime of processes)/2);
@@ 3.2 Analysis
[…]
@@ I. Balanced Selection of Jobs
In equation 1 the ratio tracks both how much a process has been waiting before execution, and how much progress has been made, before scheduling. In the numerator, (bursttime - remainingtime) is the measure of the amount of bursttime which has been consumed till a particular iteration, or in other words, this is the progress3.2 of a process till the current iteration. The waitingtime is a measure of the total amount of time one process has waited till the time it gets under execution.
@@ III. Limiting Preemption using Fairness ratio
    Assume that there are no processes in the queue ($|P|=0$, $P$: number of processes in the queue).
@@
Between 2 consecutive iterations, there is at max 1 preemption
@@
Calculating the Waiting & Response Time
@@
2. Next, if $k$ processes are selected from a total of $n$ processes, this implies that $n-k$ processes remain unexecuted. Each of these $n-k$ processes (starting from $(k+1)^{\text{th}}$ to $n^{\text{th}}$process) waits for the entire time quantum $tq$. Thus,
@@
Where: - $rt(p_{i})$ is the remaining time for process $p_{i}$. - $tq$ is the time quantum.
@@
reason: As mean and median for a single element is equal so the the timeQuantum = remainingtime for the job, as a result, it completely executes the job.
The timeQuantum is always greater than zero if there is at least one process in the batch from proposition 3.1. Now, the algorithm ’chooses’ jobs based on their latest fairnessRatio, sorts the whole batch based on the same, and calculates the timeQuantum. From the very front of the batch, it executes the processes one by one completely without any preemption. However, the very last process of an iteration may not get executed completely if the leftover time quantum is less than the current remaining time of the process.
Eventually, out of the $k$ ’chosen’ processes from the front of the batch, at least $k-1$ processes get executed fully, and at most one job is left to ’completely’ get executed. This very process will come up in subsequent groups for leftover execution. Therefore, at most one job from the ’chosen’ set of jobs at each iteration will appear in its subsequent iterations, while the rest of the jobs would be completely executed.
So, the maximum number of remaining jobs after the $k$-th iteration is:
@@
with a trivial base Case and remark 1:
@@
Where: - $R_{k}$ represents the maximum number of remaining jobs after the $k$-th iteration. - $p_{k}$ is the number of jobs selected in the $k$-th iteration.
@@
The exact number of jobs, didn’t get response after the $k$-th iteration is calculated as:
@@
Where: - $R_{k}$ represents the maximum number of remaining jobs after the $k$-th iteration. - $p_{k}$ is the number of jobs selected in the $k$-th iteration.
Importance of a Suitable Time Quantum
In a queue of processes, even when executing just one process for the minimal feasible time, (say as ”t”) for the sake of achieving utmost fairness, we must re-calibrate the fairness ratio of every process within the batch after each ”t” unit of time. So roughly, we can say after t unit of time, the dynamics of the whole batch change, prompting us to decide on the next job to execute. Given that we are operating within a uni-processing environment, at any given moment, only a single job is active. It is important to note that ”t” can vary significantly, with the potential to be arbitrarily small or large. If ”t” is exceedingly small, process progress becomes exceedingly nominal, while an arbitrarily large ”t” renders the algorithm’s behaviour akin to First-Come-First-Serve (FCFS) for the majority of the time. In traditional round-robin scheduling, ”t” or the time slice is often selected empirically, giving rise to issues concerning process progress. In our algorithm, as we look at fairness, the ideal approach is to run the algorithm after an arbitrarily small ”t”. Nevertheless, it is clear that this is practically unfeasible due to exuberant computation, preemption, and arbitrarily small progress. Which contradicts our objective of balancing fairness and efficiency. This is why we have diligently explored various formulas across current literatures [27, 28, 29, 30, 31, 32], and have finalised, inspired by the empirical work presented by the authors in [33].
3.3 Analysis of the Batch: Best and Worst Cases
Here our focus would be on best and worst case for average waiting time and average response time for the batch. This primarily means we’d be investigating if there exists any distribution under which one algorithm gives minimal waiting and/or response time, similarly for maximum waiting and/or response time. In the general case, these analyses are distribution and algorithm-dependent but in order to give a conclusive overview of where, in terms of, both average waiting and response time, our algorithm is comparatively standing.
Minimal Average Response Time
We’re considering this as the best case for response time. Under our current setting, out of n jobs in a batch, only 1 can be under execution at a time and rest ’n-1’ job wait till there is any context switching. The greedy way is allocating each task the minimum possible execution time, regardless of its burst time and moving on to the next processes. For example, if there are ’n’ number of tasks and if each task is granted k unit of time for the same, the first process experiences no initial delay, the second incurs a k-unit wait due to the first task, and so on, with the n-th process waiting for k unit for each preceding task. In this scenario, the average response time is calculated as:
$\frac{0+k+2k+\ldots+k(n-1)}{n}=\frac{k(n-1)}{2}$
remark: Generally, we take k = 1 in normal settings.
Minimal Average Waiting Time
This is the best case for waiting time. It is a well-known fact that the average waiting time is minimised by the SPT or SRTF Rule. [3, 34]

### v4

# Revitalising the Single Batch Environment: A ‘Quest’ to Achieve Fairness and Efficiency

@@ Abstract
In the realm of computer systems, efficient utilisation of the CPU (Central Processing Unit) has always been a paramount concern. Researchers and engineers have long sought ways to optimise process execution on the CPU, leading to the emergence of CPU scheduling as a field of study. In this research, we have analysed the single offline batch processing and investigated other sophisticated paradigms such as time-sharing operating systems and wildly used algorithms, and their shortcomings. Our work is directed towards two fundamental aspects of scheduling: efficiency and fairness. We propose a novel algorithm for batch processing that operates on a preemptive model, dynamically assigning priorities based on a robust ratio, employing a dynamic time slice, and utilising periodic sorting to achieve fairness. By engineering this responsive and fair model, the proposed algorithm strikes a delicate balance between efficiency and fairness, providing an optimised solution for batch scheduling while ensuring system responsiveness.
@@ 1 Introduction
00footnotetext: The authors, hereby, acknowledge that a preprint for the same manuscript can be found here: https://export.arxiv.org/pdf/2308.10062.
@@
The roots of CPU scheduling [1], [2] can be traced back to the early days of computing when batch processing systems were prevalent. In these systems, a set of fixed-size jobs was submitted as a batch to the computer, and the CPU had to execute them one after another. This approach suffered from inefficiencies as the CPU lacked responsiveness or processes waiting for the completion of others. Subsequently, to overcome certain limitations, scheduling algorithms were devised to manage the execution order of jobs, aiming to minimise idle time and maximise CPU utilisation.
Efficient scheduling not only impacts system performance but also has significant economic implications. In today’s digital era, where computational power is a valuable resource, optimising CPU scheduling can lead to substantial cost savings. By ensuring high throughput of the CPUs, organisations can complete computational tasks faster, reduce energy consumption, and ultimately save resources.
In addition to efficiency, the consideration of fairness in job selection is a crucial aspect of schedulers that is frequently under-addressed [3]. However, fairness is often viewed as a subjective metric lacking a universally agreed-upon definition in various task-scheduling contexts [4].
@@
In this study, we return to the fundamental scheduling paradigm of batch processing. Here, we have a single, fixed-size queue of jobs, each with predetermined burst times. A single machine processes the queue, with preemption allowed between jobs. Despite the growing demand for advanced scheduling algorithms in recent times, this very foundational paradigm has comparatively been overlooked. As a result, the potential usage of the paradigm has also not been utilised across various domains [5]. We acknowledge that to enhance the applicability of this paradigm in today’s computing environment, we need to design algorithms that are not only succinctly efficient but also sufficiently fair [6, 7, 8]. As most of the commonly used algorithms are primarily for time-sharing and multiprocessing systems [1], the question of what feasible means to measure fairness and efficiency for this setting has not been effectively addressed. Our work addresses this question by first analysing these measures and then analyzing the classical algorithms commonly used in time-sharing and multi-programming systems [1] with respect to these measures to understand how they perform in terms of both efficient and fair distribution of job selections across a diverse set of job clusters. Lastly, through our efforts, we propose a novel algorithm that achieves both efficiency and fairness and outperforms traditional algorithms in striking a balance between these two factors. Our algorithm strives to enhance system productivity, reduce costs, and improve the overall user experience in various batch or batch-like computing environments. Specifically, our contributions are summarized as:
  • We introduced a novel algorithm concerning faithfulness and efficiency in uni-processing batch environment;
  • We theoretically analysed and experimentally demonstrated its efficiency under a diverse and robust experimental setup and compared with other wildly used policies in the light of balancing efficiency and fairness;
  • We furthermore come up with an optimised version of our algorithm to reduce additional computational overhead.
Across the manuscript, we’ve interchangeably used the terms ‘Job’, ‘Task’, and ‘Process’.
@@ 2.1 Glossary
Over time, researchers have proposed enhancements and improvements to classical algorithms (see Appendix A Appendix A). For instance, variations of Round Robin, such as variations of Weighted Round Robin (WRR) [9], [10] and Multilevel Queue Scheduling, have been introduced to address the limitations of strict time slicing. Additionally, policies like Priority Scheduling and Multilevel Feedback Queue Scheduling have been developed to incorporate process priority and dynamically adjust scheduling parameters. Recent advancements in CPU scheduling have focused on adaptive and intelligent algorithms, leveraging machine learning and optimisation techniques [11], [12], [13]. These algorithms aim to dynamically adapt to workload patterns, predict burst times, and improve system performance. Examples include Reinforcement Learning-based scheduling [14], [15] algorithms, fuzzy-logic based heuristics [16], [17], [18] and Genetic Algorithm-based approaches [19]. While these advanced frameworks for complex operating systems take huge additional computations, other strategies, like backfilling algorithms [20], fair-share schedulers [21], gang scheduling [22], deadline-based schedulers [23] etc are complex to implement, cost computational overhead and do not comprehend the inter-dependencies of different attributes in an easy manner and not always suitable and/or applicable to the batch processing. The wildly used algorithms in scheduling has been discussed in Appendix A (Appendix A).
@@
A batch processing system can be visualised as a fixed-size array of jobs. In this research, we have considered the case where the jobs have their predetermined bursttimes. This is the most fundamental paradigm in scheduling theory which along with its several variants serves as the cornerstone for some of the most important industrial processes as well as in various domains in data-engineering [8], big data and high-performance computing [5] etc. Now, our ‘Quest’ is to analyse this very simple batch paradigm with the intention to investigate, primarily the behaviour of the two metrics researchers are mostly interested in: Efficiency and Fairness.
@@ I. Efficiency
Batch processing is heavily used in industrial setups [24, 25] where we are mostly concerned with the completion of all the jobs. In our work, there is no explicit deadline, penalty or priority of jobs. This is why, our primary means to quantify the efficiency is to measure the average waiting and turnaround time for the whole batch. Researchers, in the past, had also considered the number of tardy jobs [26], maximum lateness [27], and sometimes makespans [28] in a set of derivatives of the same paradigms which is not suitable and/or applicable for our current objectives of this work.
@@ II. Fairness
Efficiency in queuing systems generally has standard metric(s) and definitions to measure the performance of policies. Fairness, unlike efficiency, doesn’t have a universally accepted metric to deal with [4]. Experts had mostly related fairness in terms of selection and providing proportionate timeQuantum to jobs [3, 4] in the past. However, to the best of our knowledge, there is no universal quantifiable metric of fairness for single-batch processing in a uni-processing system and the aforesaid improvements are also not primarily intended for fairness in our setting.
In our study, our primary focus is on comparing algorithms based on their efficiency, i.e. in terms of minimizing average turnaround and waiting times as stated above. However, efficiency is not the sole consideration, as practitioners also emphasize the system’s average response time [29], which is the duration from a job’s arrival to its first response. In our context, where all jobs are present in the batch and no new jobs are added during processing, depending on the selection of the jobs at each iteration responsetime provides an overview of how fairly jobs have been selected. This is why, we quantify fairness by calculating the average responsetime for the entire batch. For our algorithm, at each iteration, we calculate response time for all eligible jobs and then compute the mean of these responsetimes when no jobs are left, generating the average response time for the whole batch.
To summarise, we evaluate algorithms using three key parameters: average waitingtime, average turnaroundtime, and average responsetime. The average waitingtime and average turnaroundtime are indicators of efficiency, while the average responsetime reflects the fairness of the algorithms. A competitive algorithm not only excels in efficiency but also maintains a maintains a satisfactory level of fairness. Lower turnaround and waiting times indicate higher efficiency, while a lower average responsetime indicates fairer job selection.
3 Proposed Work
@@
While empirically going through examples of process clusters, a very thoughtful observation for us was to discover the fact that any cluster of processes is never discrete throughout execution. In other words, the current attributes of a particular cluster of processes will inevitably change after some interval and we would be dealing with different types of sub-problems (here cluster of processes) in each cycle. In a multilevel queue or multilevel feedback queue for example, if we have an interval of ‘t’ units and a set of algorithms ‘S’, we may halt after each ‘t’ unit of time, analyse the system, and can choose the ‘most suitable’ algorithm from ‘S’. For a tiebreaker amongst comparable algorithms, we can even take their space and time complexity or other parameters into account. Multilevel queues and multilevel feedback queues (MLQs/MLFQs) are computationally intensive [1] and can be overkill when striking a balance between efficiency and fairness, this very observation was the main inspiration to come up with a metric that can address the crucial aspects and intricacies of the process attributes after a certain period. Our approach is primarily experimental and substantially empirical in nature. After experimenting with several ways of measuring the changes in various parameters and their interdependencies, we have finalised the ratio in the algorithm below based on a few aspects elaborated in 3.1 section. While this may not be as sophisticated as MLQ / MLFQ, the interdependencies of attributes, if addressed effectively, lead to a near-optimal solution. With the potential to address the limitations of traditional batch processing, our approach opens doors to more conclusive approaches in this domain.
@@
[…]
@@ 3.1 Analysis
$\textit{fairnessRatio}=\frac{bursttime-remainingtime+\textit{waitingtime}}{bursttime*preemptioncount}$ (1)
@@ I. Balanced Selection of Jobs
In equation 1 the ratio tracks both how much a process has been waiting before execution, and how much progress has been made, before scheduling. In the numerator, (bursttime - remainingtime) is the measure of the amount of bursttime which has been consumed till a particular iteration, or in other words, this is the progress3.1 of a process till the current iteration. The waitingtime is a measure of the total amount of time one process has waited till the time it gets under execution.
@@ III. Limiting Preemption using Fairness ratio
    Assume that there are no processes in the queue ($|P|=0$, $P$: set of processes in the queue).
@@
Between 2 consecutive iterations, there is at max 1 preemption.
@@
3.2 Formulation of the Waiting and Response Time
@@
2. Next, if $k$ processes (say from 1 to $k^{\text{th}}$ processes) are selected from a total of $n$ processes, this implies that $n-k$ processes remain unexecuted. Each of these $n-k$ processes (starting from $(k+1)^{\text{th}}$ to $n^{\text{th}}$process) waits for the entire time quantum $tq$. Thus,
@@
Where: $rt(p_{i})$ is the remainingtime for process $p_{i}$. $tq$ is the timeQuantum.
@@
reason: As the mean and median for a single element are equal so the the timeQuantum = remainingtime for the job, as a result, it completely executes the job.
The timeQuantum is always greater than zero if there is at least one process in the batch from proposition 3.1. FairBatch starts executing the processes one by one completely without any preemption however, the very last process of an iteration may not get executed completely if the leftover time quantum is less than the current remaining time of the process. As a result, out of the $k$ ‘chosen’ processes, at least $k-1$ processes get executed fully, and at most one job is left to ‘completely’ get executed. This very process will come up in subsequent groups for leftover iterations. Therefore, at most one job from the ‘chosen’ set of jobs at each iteration will appear in its subsequent iterations, while the rest of the jobs would be completely executed.
So, the maximum number of remaining jobs after the $k^{\text{th}}$ iteration is:
@@
with a trivial base case, conditioned to remark 1:
@@
Where: - $R_{k}$ represents the maximum number of remaining jobs after the $k^{\text{th}}$ iteration. - $p_{k}$ is the number of jobs selected in the $k^{\text{th}}$ iteration.
@@
The exact number of jobs, that didn’t get a response after the $k^{\text{th}}$ iteration is calculated as:
@@
Where: - $R_{k}$ represents the maximum number of remaining jobs after the $k^{\text{th}}$ iteration. - $p_{k}$ is the number of jobs selected in the $k^{\text{th}}$ iteration.
remark: Since the turnaroundtime is just the sum of the bursttime and waitingtime, and the bursttime is constant in our setting, our formulations focuses on the variable parameters, which are waitingtime and responsetime.
3.3 Further Analysis of the Parameters:
In our investigation, we will analyze when our setting achieves the minimum and maximum waiting and response time.
  • Minimal Average Response Time:
    Under our current setting, out of n jobs in a batch, only 1 can be under execution at a time and rest ‘n-1’ job wait till there is any context switching. The greedy way is allocating each task the minimum possible execution time, regardless of its bursttime and moving on to the next processes. For example, if there are ‘n’ number of tasks and if each task is granted ‘k’ unit of time, the first process experiences no initial delay, the second incurs a k-unit wait due to the first task, and so on, with the n-th process waiting for k unit for each preceding task. In this scenario, the average response time is calculated as:
    $\frac{0+k+2k+\ldots+k(n-1)}{n}=\frac{k.n(n-1)}{2}$
    remark: Usually under general settings, k = 1 [1].
  • Minimal Average Waiting Time:
    It is well-known that the SJF/SRTF (preemptive version of SJF) Rule minimises the average waiting time. [30, 31]